# AnomalyMatch: Discovering Rare Objects of Interest with Semi-supervised and Active Learning

### v1

@@ 3.5 Implementation and Scalability
Our pipeline is built to ensure scalability, robustness, and reproducibility. Initially derived and extensively refactored from an existing PyTorch implementation of FixMatch [^1] https://github.com/LeeDoYup/FixMatch-pytorch — PyTorch code for FixMatch (Accessed 2025-03-28) , our architecture has been optimized for large-scale datasets, supporting efficient handling of datasets on terabyte scale (e.g. 100 million Hubble images) [10]. We handle data in HDF5 format to optimise performance during training and inference [27].
Practically important hyperparameters include the number of unlabelled images per training iteration (limited by GPU memory) and image input sizes, which directly influence training efficiency. Supported input formats include single-channel and RGB imagery (PNG, TIFF, JPEG).
Continuous integration and continuous deployment workflows incorporating over 70 unit tests ensure code robustness, reliability, and maintainability. This is essential given the complex interaction of semi-supervised and active learning components. Pending ESA licensing processes, the complete source code will be publicly available on GitHub  [^2] https://github.com/esa/AnomalyMatch, encouraging community collaboration and ongoing development.
@@ 4 Results
4.1 Datasets and Experimental Setup
We evaluate AnomalyMatch on two diverse datasets, one related to our primary application area in astronomy (GalaxyMNIST), and one standard benchmark commonly used in machine learning research (miniImageNet). MiniImageNet was specifically chosen as it is computationally more tractable than the full ImageNet dataset, making it well-suited for our experimentation.
GalaxyMNIST
[28] comprises annotated greyscale galaxy images derived from the Galaxy Zoo citizen science project [8]. Images originally have a resolution of $64\times 64$ pixels, which we upsampled to $224\times 224$ pixels for consistency across experiments. The dataset contains four morphological classes annotated by volunteers: Smooth Round, Smooth Cigar-shaped, Edge-on Disk, and Unbarred Spiral. For evaluation, we treated each morphological type sequentially as anomalies, using the other three classes as nominal samples.
miniImageNet
[29] is a widely used benchmark in few-shot learning. It consists of colour images grouped into 100 classes and is taken from the larger ImageNet dataset. Each image was resized from its original resolution to $224\times 224$ pixels. We deliberately selected five distinct object classes as anomalies to evaluate our model’s performance across diverse visual features: piano, hourglass, guitar, printer, and orange.
Figure 3 shows representative examples from both datasets.
@@ 4.1 Datasets and Experimental Setup
Figure 3: Sample images from GalaxyMNIST (top) and MiniImageNet (bottom). Red borders indicate anomaly classes used during evaluation, while nominal samples are shown with white borders. For GalaxyMNIST all classes were tested as anomaly class in a dedicated run.
For GalaxyMNIST, we started experiments with a small labelled set of 40 images (10 anomalies, 30 nominal). For miniImageNet, initial labelled data included 500 images (five anomalies, 495 nominal). In both cases, each training step utilised an additional randomly sampled pool of 10,000 unlabelled images.
To simulate active learning, we performed three consecutive training cycles of 100 iterations each. After each cycle, the 10 highest-scoring anomalies and 10 highest-scoring false positives (nominal images with high anomaly classification scores by the model) were labelled and incorporated into the training set for the next cycle. The models are finally evaluated on all unlabelled images.
@@
The FixMatch algorithm [5] formed the basis of our semi-supervised training procedure. Key hyperparameters included an exponential moving average (EMA) momentum of 0.99, confidence threshold ($\tau$) of 0.95, temperature of 0.5, unsupervised-to-supervised loss ratio of 1.0, a batch size of 16, learning rate of 0.0075 (with SGD optimizer, momentum 0.9), and weight decay of $7.5\times 10^{-4}$. We followed empirical results from [16, 25] in the selection of parameters.

### v2

@@ 2.5 Implementation and Scalability
Our pipeline is built to ensure scalability, robustness, and reproducibility. Initially derived and extensively refactored from an existing PyTorch implementation of FixMatch [^2] https://github.com/LeeDoYup/FixMatch-pytorch – PyTorch code for FixMatch (Accessed 2025-03-28), our architecture has been optimised for large-scale datasets, supporting efficient handling of datasets on terabyte scale (e.g. 100 million Hubble images O’Ryan & Gómez, 2025). We handle data in HDF5 or Zarr format to optimise performance during training and inference (Folk et al., 2011).
Important hyperparameters include the number of unlabelled images per training cycle (limited by GPU memory) and image input sizes, which directly influence training efficiency. Supported input formats include single-channel and RGB imagery (PNG, TIFF, JPEG) and FITS files. In the case of FITS files AnomalyMatch can specify one or 3 extensions to use. Furthermore, the user can select between different normalisation options such as the simple conversion, logarithmic or Asinh stretching or a Z-Scale based linear stretch.
Continuous integration and deployment workflows incorporating over 70 unit tests ensure code robustness, reliability, and maintainability. This is essential given the complex interaction of semi-supervised and active learning components. The complete source code of AnomalyMatch is publicly available on GitHub [^3] https://github.com/esa/AnomalyMatch, encouraging community collaboration and ongoing development.
@@ 3 Results
3.1 Experimental Setup
We evaluate AnomalyMatch on three datasets: miniImageNet (Vinyals et al., 2016), a common benchmark used in machine learning research; GalaxyMNIST (Walmsley et al., 2022), a dataset pertinent to our primary application area, astronomy; and a subset of $\approx 60,000$ galaxy images extracted from Galaxy Zoo 2 (Willett et al., 2013) which were used for the ’Galaxy Challenge’ on Kaggle [^4] https://www.kaggle.com/c/galaxy-zoo-the-galaxy-challenge. The initial configurations (label counts and anomaly ratios) varied by dataset. Detailed descriptions of each dataset, along with their respective initial configurations, are provided in the following Sections. Figure 3 shows representative examples from both datasets.
@@ 3.1 Experimental Setup
Figure 3: Sample images from GalaxyMNIST (top), Galaxy Zoo 2 (centre), and MiniImageNet (bottom). Red borders indicate anomaly classes used during evaluation, while normal samples are shown with white borders. For GalaxyMNIST all classes were tested as anomaly class in a dedicated run.
@@
The FixMatch algorithm (Sohn et al., 2020) formed the basis of our semi-supervised training procedure. Key hyperparameters included an exponential moving average (EMA) momentum of 0.99, confidence threshold ($\tau$) of 0.95, temperature of 0.5, unsupervised-to-supervised loss ratio ($\lambda$ in Equation 1) of 1.0, a batch size of 16, learning rate of 0.0075 (with a stochastic gradient descent (SGD) optimiser, momentum 0.9), and weight decay of $7.5\times 10^{-4}$. We followed empirical results from Gómez & Meoni (2021); Östman et al. (2023) in the selection of parameters.
To simulate active learning for all three datasets we performed three consecutive training cycles of 100 iterations each. At the end of each cycle, the 10 highest-scoring anomalies and 10 highest-scoring false positives (normal images with high anomaly classification scores by the model) were labelled and incorporated into the training set for the next cycle. The models were finally evaluated on all unlabelled images.